# GENIAL: Generative Design Space Exploration via Network Inversion for Low Power Algorithmic Logic Units

### v1

@@ Abstract
As AI workloads proliferate, optimizing arithmetic units is becoming increasingly important to reduce the footprint of digital systems. Conventional design flows, which often rely on manual or heuristics-based optimization, are limited in their ability to thoroughly explore the vast design space. In this paper, we introduce GENIAL, a machine learning-based framework for the automatic generation and optimization of arithmetic units, more specifically multipliers.
At the core of GENIAL is a Transformer-based surrogate model trained in two stages, involving self-supervised pretraining followed by supervised finetuning, to robustly forecast key hardware metrics such as power and area from abstracted design representations. By inverting the surrogate model, GENIAL efficiently searches for new operand encodings that directly minimize power consumption in arithmetic units for specific input data distributions. Extensive experiments on large datasets demonstrate that GENIAL is consistently more sample efficient than other methods, and converges faster towards optimized designs. This enables to deploy a high-effort logic synthesis optimization flow in the loop, improving the accuracy of the surrogate model. Notably, GENIAL automatically discovers encodings that achieve up to 18% switching activity savings within multipliers on representative AI workloads compared with the conventional two’s complement. We also demonstrate the versatility of our approach by achieving significant improvements on Finite State Machines, highlighting GENIAL’s applicability for a wide spectrum of logic functions. Together, these advances mark a significant step toward automated Quality-of-Results-optimized combinational circuit generation for digital systems.
Index Terms:
RTL Generation, Logic Synthesis, High-level Synthesis, Artificial Intelligence, Machine Learning, Design Space Exploration
@@ I Introduction
This work aims to provide a modern look at how values are represented in binary format, with the ultimate goal of reducing the energy consumption of algorithmic logic units in AI accelerators and associated communication devices. An encoding is the map between a set of values and their associated binary representations. There are numerous well-recognized encodings such as the two’s complement and sign-magnitude for fixed points numbers[3], the IEEE standard floating point representations[1], and the more moderns brain float from Google[16, 41] and HiFloat from Huawei[23]. Most of these encodings were designed to support fast and efficient binary arithmetic operations, but not specifically to reduce internal switching activity, which is the main source of power consumption in compute-bound applications. The broadly implemented encoding for signed integer values, the Two’s Complement (TC), exhibits significant bit-level toggling for operands values close to zero, and may therefore be suboptimal in terms of energy efficiency for workloads with value distributions centered around zero[3]. The sign-magnitude encoding demonstrates less toggling around zero and, as a result, shows reduced switching activity for the multiplication operations. This raises the question, are there encodings that could minimize the switching activity even further than sign-magnitude.
@@
The proliferation ofLarge Language Models (LLM)-based chatbots and other compute-intensive generative applications is continuously increasing the global energy demand of data centers. On top of being hosted on cloud platforms, many LLMs are being deployed at the edge in phones, laptops, etc.. In both scenarios, weights and activations quantization to a reduced set of integer values is a standard solution to reduce the computational, communication and memory costs, as it allows to use low-precision binary numbers. Some work show that 4b quantization can have negligible impact on the accuracy of such models[5]. On top of that, those workloads tend to rely on numerous matrix multiplications, whose values are mostly concentrated around zero[12]. Under such considerations, the potential of finding better encodings for energy efficient low-precision computations has become very timely.
@@
To achieve this, we propose GENIAL: a framework that enables fast, yet powerful, binary encoding discovery for optimizing metrics such as power or area of encoding-dependent circuits. While we mainly focus on multiplier units, we also show that our framework is suitable for other circuits. This work brings forth the following contributions:
@@
  • We introduce GENIAL, a machine learning-based Design Space Exploration (DSE) framework for the discovery of novel encodings optimized to reduce power consumption of an arithmetic circuit.
@@
  • We perform comprehensive ablation studies to justify each component of the framework.
@@
  • We show that generation by Network Inversion drastically accelerates the DSE by demonsrating superior sample efficiency.
@@
  • We thus deploy a high effort Logic Synthesis Optimization (LSO) flow to maximize the accuracy of the surrogate model inherent to GENIAL.
@@
  • We execute a DSE at scale for 4b multipliers, where GENIAL reaches the Pareto front for Switching Activity (SwAct) and area of multipliers, allowing to discover a new 4b binary encoding whose energy efficiency is very close to the one of the Sign-Magnitude Extended (SME).
@@ II-A Multiplier Design and Data Formats
Research on multiplier design has long centered on minimizing transistor count, reducing power consumption, and improving timing characteristics. Techniques such as Booth encoding [6] and the Wallace tree [40] were among the earliest to address efficient partial product handling, and remain influential in modern designs. Over time, these ideas have evolved into higher performance variants, particularly through modifications to Booth’s original method [45]. The choice of numerical representation also has a direct impact on SwAct, which in turn affects dynamic power. TC is widely used for signed fixed-point arithmetic, but its carry chains can lead to increased bit transitions and higher dynamic power usage [18]. The sign-magnitude encoding, by comparison, mitigates this by toggling fewer bits when values are near zero [39]. Beyond these standard formats, custom encodings tailored to arithmetic circuits have also been proposed to improve power efficiency in multipliers [9].
@@ II-B Heuristic and Bayesian Optimization
Early efforts design space exploration and optimization relied on exhaustive enumeration and meta‐heuristics such as simulated annealing to explore small design spaces. For instance, the binary switching algorithm, a method for optimizing mappings[46], was applied to discover encodings that reduce bit error rates in communication systems [32]. More principled explorations come from Bayesian optimization (BO) with Gaussian‐process surrogates; later, deep neural networks were introduced to scale BO to higher dimensions [33]. The BANANAS framework further improved convergence by combining BO with a path‐based neural predictor [42].
II-C Reinforcement Learning, VAE, and Diffusion
Deep reinforcement learning has been used to select effective sequences of logic‐synthesis transformations[15], and multi‐agent RL has been applied to multiplier designs[11, 47]. Variational autoencoders were also deployed for circuit interpolation[35].
II-D Other Hardware Optimization Methods
In computer vision, deep generator networks synthesize inputs that maximally activate neurons [25], and Plug & Play Generative Networks perform iterative latent‐space optimization [24]. Feature‐visualization techniques decode preferred stimuli [27]. In protein design, inversion of AlphaFold’s structure predictor enables state-of-the-art protein sequence generation [17, 13]. These model‐inversion methods demonstrating powerful generation capabilities greatly inspired GENIAL’s surrogate‐inversion for combinational logic.
Another notable solution is Diffusion[14], a method originally used for image generation[10, 14] and black‐box optimization[19]. Diffusion for hardware DSE has already been tested adapted[29] and show promising capabilities for DSE. Finally, LLM and agents based optimization is also recently showing an rising interest. FunSearch uses transformer agents to guide code‐generation tasks[31], and AlphaEvolve employs LLMs with evolutionary strategies to propose design transformations[26]. Similarly, [44] optimizes Register-Transfer-Level (RTL) code. Instead of text, Circuit Transformer models directly learn on representation of the data structures to be optimized [21, 22]. Although extremely promising, LLM‐based methods for low‐level hardware synthesis remain nascent and can be very costly in their execution.
@@ III-A Overview
GENIAL is a framework aimed at exploring the space of operand encodings to minimize specific Quality of Results (QoR) metrics of a circuit—such as area or power. As illustrated in Fig. 1, it consists in an iterative DSE process built with the following components:
@@
At each iteration, new designs (also called prototypes) are suggested, and further processed with EDA tools. Relevant QoR metrics are extracted, and the QMP is finetuned on the entire dataset acquired so far. In this work, the term ”design” is used to refer to all logic circuits that have the same truth table, i.e., the same binary logic function. The terms ”circuits” or ”graph” are used to refer to the different versions of a same design.
Figure 1: GENIAL framework overview
@@ III-B1 Definition of Encoding
An encoding is the mapping between a set of values and their associated binary representation. E.g., in the TC encoding, the binary representation of the value 3 is 0011 on 4 bits. Encodings can be represented as maps of type $\text{value}_{0}$:$\text{representation}_{0}$, …} or as binary 2D tensors (Fig. 2). There are $(2^{4})!\approx 2.1\cdot 10^{13}$ 4-bit binary encodings, which defines the size of the search space for the 4b multiplier scenario.
@@ III-B2 Abstracted Design Representation
To interface with a machine learning pipeline, we must derive a representation of a design that can serve as input to a neural network. Since our focus is on exploring different encodings that implement the same logic function (binary multiplication), we represent each design solely by its encoding, formatted as a tensor. The columns correspond to the ordered individual bits of the binary representation, while the rows correspond to the ordered values being encoded. The TC and SME encoding representations are visible Fig. 2.
@@ III-B3 RTL Design Description
Design generation is challenging as the RTL must adhere to a set of rules to be synthesizable. To simplify this process, we represent each design as a Look-Up-Table (LUT) circuit, delegating all optimizations to the synthesis tool. Each design is generated by concatenating together the binary representations of the input operands into a selection vector, and assigning the output as the result of the corresponding operation. To ease integration with existing systems, the multiplier output is always in TC, and we explore only custom input encodings. For instance, a 4-bit two-input and 8-bit single-output integer multiplier would be represented as follows in Verilog, assuming the TC encoding in input:
@@ III-C EDA Task Launcher
The task launcher performs a sequence of EDA tasks on each generated design, including synthesis, post-synthesis simulation, power estimation, and finally QoR metric extraction. Leveraging different tools and flow scripts enables the extraction of various QoR metrics that capture key characteristics of a design. These metrics include Number of Transistors (NT), post-synthesis simulation-based Switching Activity or Power (P), area footprint, number of Major-Inverter-Gate (MIG) nodes (NMIG), and others. In the following, we present the different flows used in this work, which are also summarized Table I.
III-C1 Espresso-based Flow (ESPRS)
Espresso [7] is a widely used, efficient logic minimization algorithm that reduces the size of a design’s truth table. After minimization, we compute the complexity metric (CPX) by summing the counts of AND gates (in minterms) and OR gates (in maxterms).
@@ III-C2 OpenRoad Flow (OPENR)
The OpenRoad Flow Scripts[2] are a fully open-source EDA framework that allow automated synthesis and place and route operations. Internally, it synthesizes the design using Yosys [43] and performs place and route of the design using OpenROAD [36]. Combined with the OpenSTA [37] tool, this setup enables the extraction of delay and power information from the synthesized designs. The primary metrics of interests in this flow are the Number of Transistors and Power, with the former being faster to compute, as it does not require simulation. NT is estimated using Yosys’ internal transistor count model. P is obtained by mapping the design to the ASAP7 predictive Process Design Kit (PDK)[8], followed by simulation using the open source Verilator[34] Verilog simulator (v0.53).
@@ III-C3 Custom Synthesis Flow (FLOWY)
Finally, we introduce a custom synthesis flow, highly inspired by the ones of [20, 4]. It consists in an initial synthesis with Yosys, succeeded by a high effort LSO using the Mockturtle (MT) [30] tool by iterative random walk, as illustrated Fig. 3. For each generated LUT, the high effort LSO runs for 3000 steps on 12 parallel sequences of steps, called chains. At each step, a recipe is randomly selected from 30 available options and applied once if it is a decompression script, or three times if it is a compression script. At the end of the 3000 steps, the circuit with lowest number of MIG gates is selected from each chain. These 12 circuits are then simulated with Verilator, and the one exhibiting the lowest SwAct is chosen as the final circuit for the LUT.
Figure 3: Custom high effort logic synthesis optimization flow
@@
III-C4 Switching Activity Model
To enable fast and large-scale evaluation of millions of designs, we introduce a lightweight power estimation protocol based on SwAct extraction, described below.
  1. 1.
    Stimulus generation: sample an input a 20k-long sequence of input value pairs from a normal distribution, and apply one pair of values to the multiplier’s input ports per clock cycle.
  2. 2.

### v2

@@ Abstract
As AI workloads proliferate, optimizing arithmetic units is becoming increasingly important for reducing the footprint of digital systems. Conventional design flows, which often rely on manual or heuristic-based optimization, are limited in their ability to thoroughly explore the vast design space. In this paper, we introduce GENIAL, a machine learning-based framework for the automatic generation and optimization of arithmetic units, with a focus on multipliers.
At the core of GENIAL is a Transformer-based surrogate model trained in two stages, involving self-supervised pretraining followed by supervised finetuning, to robustly forecast key hardware metrics such as power and area from abstracted design representations. By inverting the surrogate model, GENIAL efficiently searches for new operand encodings that directly minimize power consumption in arithmetic units for specific input data distributions. Extensive experiments on large datasets demonstrate that GENIAL is consistently more sample efficient than other methods, and converges faster towards optimized designs. This enables deployment of a high-effort logic synthesis optimization flow in the loop, improving the accuracy of the surrogate model. Notably, GENIAL automatically discovers encodings that achieve up to 18% switching activity savings within multipliers on representative AI workloads compared with the conventional two’s complement. We also demonstrate the versatility of our approach by achieving significant improvements on Finite State Machines, highlighting GENIAL’s applicability for a wide spectrum of logic functions. Together, these advances mark a significant step toward automated Quality-of-Results-optimized combinational circuit generation for digital systems.
@@ I Introduction
This work aims to provide a modern look at how values are represented in binary format, with the ultimate goal of reducing the energy consumption of algorithmic logic units in AI accelerators and associated communication devices. An encoding is a mapping between a set of values and their corresponding binary representations. There are numerous well-recognized encodings such as the two’s complement and sign-magnitude for fixed-point numbers[arnold2025explicitsignmagnitudeencoders], the IEEE standard floating-point representations[ieee754], and the more modern Brain Float from Google[intel-bfloat16-2020, wang2019bfloat16] and HiFloat from Huawei[luo2024ascendhifloat8formatdeep]. Most of these encodings were designed to support fast and efficient binary arithmetic operations. However, they were not specifically designed to reduce internal switching activity, which is the main source of power consumption in compute-bound applications. The broadly implemented encoding for signed integer values, the Two’s Complement (TC), exhibits significant bit-level toggling for operand values close to zero and may therefore be suboptimal in terms of energy efficiency for workloads with value distributions centered around zero. The Sign-Magnitude (SM) encoding demonstrates less toggling around zero and, as a result, shows reduced switching activity for the multiplication operations[arnold2025explicitsignmagnitudeencoders]. This raises the question: are there encodings that could decrease switching activity even further than SM?
@@
The proliferation of Large Language Models (LLM)-based chatbots and other compute-intensive generative applications is continuously increasing the global energy demand of data centers. Besides being hosted on cloud platforms, many LLMs are deployed at the edge in phones, laptops, etc. In both scenarios, quantizing weights and activations to a smaller set of integer values is a common strategy to reduce computational, communication, and memory overheads, enabling more efficient low-precision computation. Prior studies have shown that 4-bit quantization can have a negligible impact on model accuracy [NEURIPS2024_b5b93943, muller2025sinq]. Furthermore, these workloads tend to rely on numerous matrix multiplications, whose values are mostly concentrated around zero[Gongyo_2024_ACCV]. Consequently, the potential of finding better encodings for energy-efficient low-precision computations has become increasingly important.
Unlike prior approaches, such as Bayesian Optimization [grosnit2022boils] or reinforcement learning [hosny2020drills], that optimize transformation sequences or parameterized Register-Transfer Level (RTL) templates, GENIAL directly targets the operand encoding itself. This is realized through a combination of (i) self-supervised pretraining to learn robust encoding representations, (ii) legality-aware network inversion losses that preserve bijectivity, and (iii) a hierarchical EDA-in-the-loop flow with options trading off execution time and fidelity.
@@
We propose GENIAL: a framework that enables fast and effective binary encoding discovery for optimizing circuit-level metrics such as power or area. While we mainly focus on multiplier units, we also show that our framework generalizes to other circuit types. This work brings forth the following contributions:
@@
  • We introduce GENIAL, a machine learning-based Design Space Exploration (DSE) framework for the discovery of novel binary encodings that optimize arithmetic circuits for power efficiency.
@@
  • We perform comprehensive ablation studies to justify each component within the framework.
@@
  • We show that generation via Network Inversion drastically accelerates DSE, achieving superior sample efficiency compared to random or heuristic search.
@@
  • We deploy a high-effort Logic Synthesis Optimization (LSO) flow to maximize the usefulness of the surrogate model inherent to GENIAL.
@@
  • We conduct large-scale DSE experiments on 4-bit multipliers, where GENIAL reaches the Pareto front for Switching Activity (SwAct) and area, and discovers a new 4-bit binary encoding whose energy efficiency closely matches that of the Sign-Magnitude Extended (SME) format.
  • We release the full GENIAL framework, including training scripts, surrogate models, and EDA integration utilities, to foster reproducibility and further research. The code is available at: https://github.com/huawei-csl/GENIAL.
@@ II-A Multiplier Design and Data Formats
Research on multiplier design has long centered on minimizing transistor count, reducing power consumption, and improving timing characteristics. Techniques such as Booth encoding [booth1951] and the Wallace tree [wallace1964] were among the earliest to address efficient partial product handling, and remain influential in modern designs. Over time, these ideas have evolved into higher performance variants, particularly through modifications to Booth’s original method [wen_chang_high_speed_booth]. The choice of numerical representation also has a direct impact on SwAct, which in turn affects dynamic power. TC is widely used for signed fixed-point arithmetic, but its carry chains can lead to increased bit transitions and higher dynamic power usage [SwitchinginmultipliersMsc]. The SM encoding, by comparison, mitigates this by toggling fewer bits when values are near zero [waeijen2018]. Beyond these standard formats, custom encodings tailored to arithmetic circuits have also been proposed to improve power efficiency in multipliers [costa2001power].
@@ II-B Heuristic and Bayesian Optimization
Early efforts in design space exploration and optimization relied on exhaustive enumeration and meta‐heuristics such as simulated annealing to explore small design spaces. For instance, the binary switching algorithm, a method for optimizing mappings[64657], was applied to discover encodings that reduce bit error rates in communication systems [schreckenbach2003optimized]. More principled explorations come from Bayesian optimization (BO) with Gaussian‐process surrogates [grosnit2022boils]; later, deep neural networks were introduced to scale BO to higher dimensions [snoek2015scalable]. The BANANAS framework further improved convergence by combining BO with a path‐based neural predictor [white2021bananas].
II-C Reinforcement Learning and VAE
Deep reinforcement learning has been used to select effective sequences of logic‐synthesis transformations[hosny2020drills], and multi‐agent RL has been applied to multiplier designs[feng2024gomarl, zuo2023rl]. Variational autoencoders were also deployed for circuit interpolation[song2024circuitvae].
II-D Network Inversion
In computer vision, deep generator models synthesize inputs that maximally activate neurons [nguyen2016synthesizing], and Plug & Play Generative Networks perform iterative latent‐space optimization [nguyen2017plug]. Feature‐visualization techniques decode preferred stimuli [olah2017feature]. In protein design, inversion of AlphaFold’s structure predictor enables state-of-the-art protein sequence generation [jumper2021highly, goverde2023novo]. Collectively, these network inversion techniques demonstrate powerful generative capabilities and directly inspired GENIAL’s surrogate-based inversion approach for combinational logic.
II-E Other Hardware Optimization Methods
Another notable solution is Diffusion[NEURIPS2020_4c5bcfec], a method originally used for image generation[dhariwal2021diffusion, NEURIPS2020_4c5bcfec] and black‐box optimization[krishnamoorthy2023diffusion]. Diffusion for hardware DSE has already been adapted [ren2025diffuse] and shows promising capabilities for DSE. Finally, LLM- and agentic-based optimization has also recently attracted increasing interest. FunSearch uses Transformer agents to guide code‐generation tasks[romera-paredes_mathematical_2024], and AlphaEvolve employs LLMs with evolutionary strategies to propose design transformations[novikov2025alphaevolve]. Similarly, [yao2024rtlrewriter] optimizes RTL code. Instead of text, Circuit Transformer models directly learn on a representation of the data structures being optimized [li2024circuit, li2024logic]. Although extremely promising, LLM‐based methods for low‐level hardware synthesis remain nascent and can be costly in their execution.
@@ III-A Overview
GENIAL is a framework aimed at exploring the space of operand encodings to minimize specific Quality of Results (QoR) metrics of a circuit, such as area or power. As illustrated in Fig. 1, it consists of an iterative DSE process built with the following components:
@@
At each iteration, new designs (also called prototypes) are suggested, and further processed with EDA tools. Relevant QoR metrics are extracted, and the QMP is finetuned on the entire dataset acquired so far. In this work, the term design is used to refer to all logic circuits that have the same truth table, i.e., the same binary logic function. The terms circuits or graph are used to refer to the different versions of the same design.
Figure 1: GENIAL framework overview: Generator → EDA Flows → Surrogate Model → Recommender
@@ III-B1 Definition of Encoding
An encoding is the mapping between a set of values and their associated binary representations. E.g., in the TC encoding, the binary representation of the value 3 is 0011 on 4 bits. A specific encoding can be specified as a dictionary $v_{i}\mapsto c_{i}$ from value to encoded representation, or as a codebook tensor (Fig. 2). There are $(2^{4})!\approx 2.1\cdot 10^{13}$ 4-bit binary encodings, which define the size of the search space for the 4-bit multiplier scenario.
@@ III-B2 Abstracted Design Representation
To interface with a machine learning pipeline, we must derive a representation of a design that can serve as input to a neural network. Since our focus is on exploring different encodings that implement the same logic function (binary multiplication), we represent each design solely by its encoding, formatted as a tensor. The columns correspond to the ordered individual bits of the binary representation, while the rows correspond to the ordered values being encoded. The TC and SME encoding representations are visible in Fig. 2.
@@ III-B3 RTL Design Description
Generating synthesizable RTL for arbitrary operand encodings is non-trivial. To simplify this process, each design is expressed as a Look-Up-Table (LUT) circuit, allowing downstream synthesis tools to handle structural optimization. The LUT abstraction provides a technology-neutral, synthesizable representation that guarantees functional correctness for any input encoding, while outputs are consistently kept in TC format to ensure compatibility with existing digital systems. Importantly, the LUT is not the final hardware description but an intermediate form that decouples encoding legality from implementation details. Optimized RTL implementations are subsequently recovered through reliable synthesis flows such as Yosys or FLOWY. For example, a 4-bit multiplier with two inputs and one 8-bit output can be represented as:
@@
As illustrated, the two 4-bit operands are concatenated into an 8-bit input pattern. All $2^{8}$ input patterns are enumerated with their corresponding 8-bit outputs.
@@ III-C EDA Task Launcher
The task launcher performs a sequence of EDA tasks on each generated design, including synthesis, power estimation, and finally QoR metric extraction. Crucially, different metrics have different trade-offs in usefulness vs. cost: for instance, SwAct provides a close proxy to dynamic power, but requires full simulation, whereas simple complexity counts can be obtained nearly instantly. To balance accuracy and scalability, we implement three complementary flows and use them hierarchically: ESPRS for ultra-fast screening with Complexity (CPX), OPENR for mid-fidelity Number of Transistors (NT)/Power (P), and FLOWY for final SwAct/NT.
III-C1 Espresso-based Flow (ESPRS, very fast/low fidelity)
Espresso [brayton_logic_1984] is a well-established logic minimization algorithm that reduces the size of a design’s truth table. After minimization, we compute the complexity metric (CPX) by summing the counts of AND gates (in minterms) and OR gates (in maxterms). Although coarse, CPX is strongly correlated with transistor count derived by Yosys in the OPENR flow (Pearson $\rho=0.71$, see Fig. 9), making it a lightweight proxy for key QoR metrics that can be used during design space exploration.
@@ III-C2 OpenROAD Flow (OPENR)
The OpenROAD Flow Scripts [ajayi_openroad_2019] provide a fully open-source pipeline for synthesis and place-and-route. Internally, Yosys [wolf_yosys_nodate] performs the logic synthesis, while OpenROAD [the_openroad_project_openroad_2025] and OpenSTA [the_openroad_project_opensta_2025] handle placement, routing, and timing/power analysis. The primary metrics of interest in this flow are the Number of Transistors (NT) and average Power (P), with the former being faster to compute, as it does not require simulation. NT is estimated using Yosys’ internal transistor count model. To extract P, the design is mapped to the ASAP7 predictive Process Design Kit (PDK)[CLARK2016105], followed by simulation using the open-source Verilator[verilator] simulator (v0.53). Finally, P is obtained using local SwAct data combined with PDK information. This flow strikes a balance between rapid evaluation and metric fidelity representative of hardware implementation.
@@ III-C3 Custom Synthesis Flow (FLOWY)
We introduce a custom synthesis flow, highly inspired by the ones of [lee_late_2024, arnold_late_2025]. It consists in an initial synthesis with Yosys, succeeded by a high effort LSO using the Mockturtle (MT) [riener_scalable_2019] tool by iterative random walk, as illustrated in Fig. 3. For each generated LUT, the high effort LSO runs for 3000 steps on 12 parallel sequences of steps, called chains. At each step, a recipe is randomly selected from 30 available options and applied once if it is a decompression script, or three times if it is a compression script. At the end of the 3000 steps, the circuit with the lowest number of Majority-Inverter-Gate (MIG) gates is selected from each chain. These 12 circuits are subsequently simulated using Verilator, and the one with the lowest SwAct is selected as the final implementation for the LUT. This process is significantly slower than ESPRS or OPENR but yields highly optimized transistor-level designs that closely approximate the output of a real EDA backend.
Figure 3: Custom high-effort logic synthesis optimization flow
@@
III-C4 Flow Selection and Runtime Considerations
To balance accuracy and scalability, we employ a hierarchical approach using the three EDA flows introduced above:
  • ESPRS (Fast Screening): Ultra-lightweight structural proxy, allowing rapid filtering of low-quality designs based on CPX.
  • OPENR (Mid-Fidelity): Logic-synthesis pass evaluating normalized NT and coarse P, used for downselection after ESPRS.
  • FLOWY (High-Fidelity): Full LSO evaluation including SwAct, applied only to finalist candidates.
This hierarchical strategy renders end-to-end search practical by evaluating only a small subset of promising designs with the more computationally expensive flow. Table II reports typical runtime per design evaluation on our hardware (8-core CPU@3.5 GHz, no GPU acceleration for synthesis).
TABLE II: Typical runtime per design evaluation for the three EDA flows used in GENIAL.
Flow Runtime / sample #CPUs
ESPRS << 0.1 s 1
OPENR $\approx$ 1 min 1
FLOWY $\approx$ 15 min 12
In practice, ESPRS evaluates the full set of candidates generated by Network Inversion, OPENR is applied to the top fraction ranked by the surrogate model, and only a small number of finalists proceed to FLOWY for accurate power-proxy estimation.
III-C5 Switching Activity Model
For scalability, we introduce a lightweight power estimation protocol based on SwAct extraction, described below.
  1. i.
    Stimulus generation: sample an input, a 20k-long sequence of input value pairs from a normal distribution, and apply one pair of values to the multiplier’s input ports per clock cycle.
  2. ii.